40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Entry  Tue Jul 2 11:21:24 2013, Erica, DailyProgress, fiber optic, mode matching snip_of_beam_profiling_6_27.JPGP1020518.JPG
    Reply  Wed Jul 3 09:46:48 2013, Erica, DailyProgress, fiber optic, mode matching 130703-095332.jpg
       Reply  Thu Jul 11 01:40:19 2013, Erica, DailyProgress, fiber optic, mode matching P1020690.JPGP1020692.JPG
          Reply  Fri Jul 12 00:51:39 2013, Erica, Notes, fiber optic, mode matching P1020705.JPGP1020703.JPG7_11_13_beam_profile.PNG7_11_13_beam_from_fixed_lens.PNG
             Reply  Tue Jul 16 10:27:47 2013, Erica, Notes, fiber optic, mode matching, beam recombination 
                Reply  Tue Jul 16 11:02:47 2013, Erica, Notes, fiber optic, mode matching, beam recombination 130805-120628.jpg
                   Reply  Wed Jul 17 11:58:47 2013, Erica, Notes, fiber optic, mode matching, beam recombination 7_16_13_determining_polarization_axis.PNGTEK0000.PNG
                      Reply  Wed Jul 17 18:29:37 2013, Erica, DailyProgress, fiber optic, reducing drift in the recombined beam 
                         Reply  Sat Jul 20 23:39:09 2013, Erica, DailyProgress, fiber optic, reducing drift in the recombined beam 
                            Reply  Tue Jul 23 10:29:12 2013, Erica, DailyProgress, fiber optic, reducing drift in the recombined beam shielded_10s_TEK0005.PNGunshielded_10s_TEK0027.PNG
                               Reply  Tue Jul 23 11:43:00 2013, Evan, DailyProgress, Electronics Equipment, Zach's op-amp recommendations 
                               Reply  Fri Jul 26 11:07:44 2013, Erica, DailyProgress, BEAT, circuit for measuring temperature fluctations on CTN table P1030056.JPGP1030057.JPGP1030059.JPGP1030056.JPG
                                  Reply  Fri Jul 26 11:22:45 2013, Erica, DailyProgress, BEAT, circuit for measuring temperature fluctations on CTN table 
                                     Reply  Fri Jul 26 11:45:28 2013, Erica, DailyProgress, BEAT, circuit for measuring temperature fluctations on CTN table 
                                        Reply  Mon Jul 29 13:30:26 2013, Erica, DailyProgress, BEAT, adding resistor, capacitor, and sockets w_o_capacitor_7_26_13.JPGcapacitor_7_26_13.JPGP1030066.JPGP1030070.JPG
                                           Reply  Tue Jul 30 12:23:41 2013, Erica, DailyProgress, BEAT, addressed noise problem, DC signal 
                                              Reply  Thu Aug 1 00:25:02 2013, Erica, DailyProgress, fiber optic, insulating foam 731_7graph.PNG731_11graph.PNGP1030318.JPG
                                                 Reply  Fri Aug 2 12:38:41 2013, tara, DailyProgress, fiber optic, insulating foam chmbrtmp.JPGfringe.JPG
                                                    Reply  Tue Aug 6 12:31:24 2013, Erica, DailyProgress, fiber optic, insulating foam P1030403.JPG
                                                       Reply  Wed Aug 7 13:27:00 2013, Erica, DailyProgress, fiber optic, more foam, calculations for convertion from voltage to frequency P1030409.JPGP1030417.JPGP1030420.JPG
                                                          Reply  Wed Aug 7 18:53:20 2013, Erica, DailyProgress, fiber optic, noise budget, Matlab notes. noise_budget.PNGP1030427.JPG
                                                             Reply  Tue Aug 13 12:39:36 2013, Erica, DailyProgress, fiber optic, rewrote calculations for converting from voltage noise to frequency noise P1030458.JPGP1030459.JPGP1030460.JPG
                                                       Reply  Tue Aug 13 12:28:20 2013, Erica, Notes, Electronics Equipment, new AD 590 circuit and its noise calculation, beam alignment P1030461.JPGP1030454.JPGP1030456.JPGP1030457.JPGAD590circuit2.PNG
                                                          Reply  Fri Aug 16 10:35:44 2013, Erica, Notes, Electronics Equipment, AD590 msrmt, conversion to freq noise 
                   Reply  Tue Jul 30 16:19:32 2013, Erica, Notes, fiber optic, setup for beam recombination new_setup.PNGbeam_recombination2.PNG
                   Reply  Mon Aug 5 12:21:00 2013, Erica, Notes, fiber optic, calculation: phase change for fiber 130805-121449.jpg
    Reply  Tue Jul 30 15:48:47 2013, Erica, DailyProgress, fiber optic, noise for GYRO 
Message ID: 1245     Entry time: Sat Jul 20 23:39:09 2013     In reply to: 1243     Reply to this: 1251
Author: Erica 
Type: DailyProgress 
Category: fiber optic 
Subject: reducing drift in the recombined beam 
July 18
Redid inventory of the lenses and printed out a new sheet because there are fused silica and BK7 lenses, which have different focal lengths for a given radius of curvature. It only has the lenses in the box. Tara had like 6 to 8 other ones but I don't know where they are.

Looked for papers that give numbers for noise added... Like before, I'm having trouble.

got login for SVN to work

Evan helped come up with other solutions to try to reduce the drift on signal from the recombined beam. We put UHV aluminum foil on the coil of fiber to reduce the acoustic noise but it didn't seem to have much effect on the drift. Acoustic sound affected the fiber because the drift seemed to speed up a little bit when we closed the cabinet door and jumped around. When the lab door was closed, there was a spike in the signal but thing settled down fairly quickly. We turned off both Hepa filters but like yesterday, it also didn't seem to do much. We elevated the coil off the table by using the black rectangular clamps to dampen possible transfer of movement from the table to the coil.
We locked the laser to the cavity to reduce the laser noise. The error and control signals were very steady compared to the drift in the beam signal.

We calculated the change in frequency over 1 period of drift is about 5 MHz, and for the fiber to change by 1 wavelength, it would take a temperature change of 16 mK. Looking at Rana's graph of for noise calculations (http://nodus.ligo.caltech.edu:8080/PSL_Lab/1205), we calculated the rms temperature noise due to ambient air at 100 mHz (because the period of the drift was on the order of 10s) was 100mK rms, which is an order of 10 larger than the calculated change in temperature.

The fiber passes by the laser, which is emitting a bit of heat, which may be a part of what is causing the drift.
ELOG V3.1.3-